OpenStudy (anonymous):

translate the phrase into a mathematical expression and then evaluate it. forty divided by ten plus three

OpenStudy (souvik):

(40/10)+3

OpenStudy (anonymous):

Forty divided by ten would be 40/10. We then need to add three to the whole thing. So we get: \[\frac{ 40 }{ 10 } + 3 = 4+3=7\]

OpenStudy (anonymous):

evaluated the expression is equal to

OpenStudy (souvik):

40/10=4

OpenStudy (souvik):

so 40/10+3=4+3=7

OpenStudy (anonymous):

the Initial balance in a checking account is $425 find the final balance resulting from the following withdrawals and deposits -$35,$23,$120,and -$260